# Chiara Arina Centre for Cosmology, Particle Physics and Phenomenology (CP3), Université catholique de Louvain, B-1348 Louvain-la-Neuve, Belgium

### v1

@@ IV.3.1 Indirect detection
For Dirac, real vector, and most complex DM models except when it is a scalar, S-wave annihilation dominates, producing a continuum of photons, positrons, neutrinos, and antiprotons. In regions where co-annihilation or bound state formation is relevant (which corresponds to a new physics spectrum featuring a mediator mass that is at most 1.2 times larger than the DM mass), the same processes that determine the relic density within the freeze-out mechanism also generate signals for indirect detection, and calculations typically yields annihilation cross sections near the canonical value of $10^{-26}~{}{\rm cm}^{3}/{\rm s}$.
@@ IV.3.2 Direct detection
$f^{(p)}_{V,\text{tree}}=\begin{cases}2\,\lambda_{V}&\text{for coupling to }u_{R}\\ \lambda_{V}&\text{for coupling to }d_{R}\\ 3\,\lambda_{V}&\text{for coupling to }(u_{L},d_{L})\end{cases}\,,\ \ f^{(n)}_{V,\text{tree}}=\begin{cases}\lambda_{V}&\text{for coupling to }u_{R}\\ 2\,\lambda_{V}&\text{for coupling to }d_{R}\\ 3\,\lambda_{V}&\text{for coupling to }(u_{L},d_{L})\end{cases}\,,$ (57)
@@
In the parameter space of the $t$-channel mediator models discussed in this work, searches for spin-independent nucleon-DM scattering in data from the experiments XENON1T [365], XENONnT [366], and LZ [367] have achieved unprecedented sensitivity, imposing constraints even on loop-suppressed interactions. For models with couplings to first-generation quarks, next-generation experiments such as DARWIN [368, 369] are expected to explore much of the currently viable parameter space, as their reach on t he spin-independent scattering cross section exceeds the neutrino background [370]. Regarding spin-dependent proton scattering, the PICO-60 experiment [371] sets strong exclusion limits for DM masses below 500 GeV, even for Majorana DM [362]. In contrast, xenon-based detectors provide the strongest limits for spin-dependent neutron interactions.
@@ IV.3.3 Early universe constraints
CMB spectral distortions also test the nature of the dark sector. These distortions, which arise from non-standard energy injection (e.g. from particle decays or annihilations) that disrupts the thermodynamic equilibrium between photons and free electrons after BBN ($z\lesssim 10^{8}$), can be probed and thus used as constraints on models [382]. The FIRAS instrument on COBE measured the CMB energy spectrum, and it found it consistent with a perfect black-body spectrum at $T_{0}=2.725\pm 0.002$ K [383, 384]. For mediator decays with lifetimes $10^{4}$ s $<\tau<10^{13}$ s, CMB spectral distortions provide stronger constraints than CMB anisotropies, though they are still weaker than those originating from BBN. Future experiments such as PIXIE or PRISM have nevertheless the potential to significantly improve the bounds [382].
Big Bang Nucleosynthesis, which occurred at $T_{\rm BBN}\sim 0.1\,$MeV, is another critical probe of dark matter. New particles can affect the primordial abundances of light nuclei by altering the Hubble rate or the entropy density of the universe. This leads to constraints on $N_{\rm eff}$, which are comparable to those originating from the CMB [385], but less stringent than the Lyman-$\alpha$ forest bounds discussed below. In addition, if the mediator has a lifetime longer than approximately $0.1$ s, its decay can induce non-thermal nuclear reactions during or after BBN, modifying standard predictions [386, 387]. Such constraints are especially relevant for superWIMP production scenarios involving mediators with lifetimes $10^{4}$ s $<\tau<10^{13}$ s that decay into hadronic final states. For example, in top-philic scenarios, these constraints help close the existing gap between constraints stemming from Lyman-$\alpha$ forest probes and the collider limits [201, 204].
@@
$M_{X}\gtrsim\begin{cases}15\,{\rm keV}\times\left(\frac{106.75}{g_{*}(T_{\rm FI})}\right)^{1/3}&\text{for freeze-in through decays},\\ 3.8\,{\rm GeV}\times\left(R_{\Gamma}^{\rm SW}/10^{-12}\right)^{-1/2}\times\left(\frac{106.75}{g_{*}(T_{\rm SW})}\right)^{1/3}&\text{for superWIMP},\\ \end{cases}$ (66)
@@ IV.4.1.1 Minimal quark-philic models in the canonical freeze-out regime
Figure 42: Constraints on minimal simplified $t$-channel DM models from cosmological and astrophysical observables, as well as from the measured $Z$-boson visible decay width. The coloured region in the $(M_{X},M_{Y}/M_{X}-1)$ plane represents scenarios that achieve $\Omega h^{2}\simeq 0.12$, with the coupling value $\lambda$ indicated by the grey-scale colour map. The left (right) panels correspond to models with self-conjugate (complex) DM, featuring scalar (top row), fermionic (middle row), and vector (bottom row) DM particles. The hatched regions indicate exclusions from gamma-ray searches (‘ID gamma rays’), cosmic-ray antiproton searches (‘ID antiprotons’), DM direct detection via spin-independent and spin-dependent interactions (‘DD SI’ and ‘DD SD,’ respectively), and $Z$-boson visible decays (‘Z decay’). The blank upper region corresponds to scenarios requiring non-perturbative couplings, while the white region at the bottom represents the CDFO regime (see section IV.4.1.2). This figure is adapted from [25], where further details can be found.
@@
[…]
Figure 43: Cosmologically viable region of the parameter space of models featuring a DM state coupling to the right-handed top quark $t_{R}$ (with $\Omega h^{2}=0.12$), and constraints shown as functions of the DM mass $M_{X}$ and the relative mass splitting $M_{Y}/M_{X}-1$. The left panel corresponds to Majorana DM, while the right panel depicts a scenario with real scalar DM. These plots are based on the results from [201] and [27], respectively. The thick black line separates the WIMP region (above) from the CDFO region (see section IV.4.1.2), and the shaded regions denote exclusions due to the various experimental and theoretical constraints discussed in the text.
@@ IV.4.2.2 Frustrated dark matter
By construction, $2\to 2$ interactions between the DM and the SM occur only at one-loop order. Consequently, for most DM masses, DM annihilation is dominated by $XX$ scattering to at least four SM particles, which effectively reduce to $t$-channel $XX$ annihilations to a pair of mediators when the mediators can be produced nearly on-shell. For perturbative Yukawa-like DM-mediator couplings $\lambda_{X}\lesssim\sqrt{4\pi}$, efficient DM annihilation requires these $t$-channel processes to be kinematically accessible; otherwise, DM tends to be overabundant under the assumption of a standard cosmological history. However, the fDM framework allows for a wide range of mediator masses and SM couplings, enabling the construction of numerous specific realisations with large experimentally viable regions in the model parameter space. The regions of the fDM parameter space that can produce a viable thermal relic are, in principle, testable via indirect searches for DM annihilation in the cosmos. While loop-suppressed $2\to 2$ annihilation processes leading to photons or a $\gamma Z$ final state are possible, the dominant experimental signature of the models is typically an apparent excess in the continuum spectra of gamma or cosmic rays originating from DM annihilation into quarks and/or charged leptons. For fDM scenarios at or below the TeV scale, the strongest constraints thus come from Fermi-LAT searches for gamma-ray production in nearby dwarf spheroidal galaxies. On the other hand, direct detection is fully loop-suppressed since there are no tree-level diagrams for fDM scattering off SM particles. Nonetheless, constraints from direct detection experiments such as XENON1T can be quite stringent. In scenarios where mediators carry weak hypercharge, spin-independent DM-nucleon scattering via off-shell photon exchange then provides the most robust constraints, that can also be interpreted as upper limits on the DM magnetic dipole moment [362]. In the absence of such processes, box diagrams involving mediators and SM fields can sometimes generate detectable direct detection signals [360] that may lead to competitive or even exceed limits from indirect detection. More conclusive quantitative statements require, however, specifying a particular fDM construction.
@@ IV.4.2.3 Composite dark matter
Figure 50: Constraints on the parameter space of a composite $t$-channel DM scenario featuring one DM state $X$ and one mediator state $Y$ that couples to the top quark with a strength $\lambda_{t}$. Results are given in the $(M_{X},M_{Y})$ plane, and the value of the coupling $\lambda_{t}$, given through the colour map, is determined in order to reproduce the observed relic density $\Omega_{CDM}h^{2}=0.1186$. The black lines correspond to isolines of constant coupling values with $\lambda_{t}=1$, $2$, $3$, $4$ and $5$.
@@ A.1 Description of the analysis
Bin 0 EM0: $200$ GeV $<E_{T}^{\mathrm{miss}}<250$ GeV IM0: $E_{T}^{\mathrm{miss}}>200$ GeV
Bin 1 EM1: $250$ GeV $<E_{T}^{\mathrm{miss}}<300$ GeV IM1: $E_{T}^{\mathrm{miss}}>250$ GeV
Bin 2 EM2: $300$ GeV $<E_{T}^{\mathrm{miss}}<350$ GeV IM2: $E_{T}^{\mathrm{miss}}>300$ GeV
Bin 3 EM3: $350$ GeV $<E_{T}^{\mathrm{miss}}<400$ GeV IM3: $E_{T}^{\mathrm{miss}}>350$ GeV
Bin 4 EM4: $400$ GeV $<E_{T}^{\mathrm{miss}}<500$ GeV IM4: $E_{T}^{\mathrm{miss}}>400$ GeV
Bin 5 EM5: $500$ GeV $<E_{T}^{\mathrm{miss}}<600$ GeV IM5: $E_{T}^{\mathrm{miss}}>500$ GeV
Bin 6 EM6: $600$ GeV $<E_{T}^{\mathrm{miss}}<700$ GeV IM6: $E_{T}^{\mathrm{miss}}>600$ GeV
Bin 7 EM7: $700$ GeV $<E_{T}^{\mathrm{miss}}<800$ GeV IM7: $E_{T}^{\mathrm{miss}}>700$ GeV
Bin 8 EM8: $800$ GeV $<E_{T}^{\mathrm{miss}}<900$ GeV IM8: $E_{T}^{\mathrm{miss}}>800$ GeV
Bin 9 EM9: $900$ GeV $<E_{T}^{\mathrm{miss}}<1000$ GeV IM9: $E_{T}^{\mathrm{miss}}>900$ GeV
Bin 10 EM10: $1000$ GeV $<E_{T}^{\mathrm{miss}}<1100$ GeV IM10: $E_{T}^{\mathrm{miss}}>1000$ GeV
Bin 11 EM11: $1100$ GeV $<E_{T}^{\mathrm{miss}}<1200$ GeV IM11: $E_{T}^{\mathrm{miss}}>1100$ GeV

### v2

@@ IV.3.1 Indirect detection
In freeze-out scenarios where DM self-annihilation dominates and proceeds via S-wave processes (as for Dirac, real vector and most complex DM models except when it is a scalar), indirect detection signals are typically expected in the form of a continuum of photons, positrons, neutrinos and antiprotons. However, in regions where co-annihilation or bound-state formation plays an important role (typically when the mediator mass is less than about 1.2 times the DM mass), the connection between the relic density and indirect detection can weaken or break down entirely. In particular, if the relic density is set by processes involving the pair-annihilation of the co-annihilating partner (such as $YY$ annihilations to a pair of SM particles), no indirect detection signal is expected today. Nonetheless, in the mass and coupling ranges considered, annihilation cross sections near the canonical value of $10^{-26}~\text{cm}^{3}/\text{s}$ still arise in considerable parts of the parameter space.
@@ IV.3.2 Direct detection
[…]
@@
In the parameter space of the $t$-channel mediator models discussed in this work, searches for spin-independent nucleon-DM scattering in data from the experiments XENON1T [371], XENONnT [372], and LZ [373] have achieved unprecedented sensitivity, imposing constraints even on loop-suppressed interactions. For models with couplings to first-generation quarks, next-generation experiments such as DARWIN [374, 375] are expected to explore much of the currently viable parameter space, as their reach on the spin-independent scattering cross section exceeds the neutrino background [376]. Regarding spin-dependent proton scattering, the PICO-60 experiment [377] sets strong exclusion limits for DM masses below 500 GeV, even for Majorana DM [368]. In contrast, xenon-based detectors provide the strongest limits for spin-dependent neutron interactions. While this review was being finalised, impressive new results were released by PandaX-4T [378] and LZ2025 [379]. In particular, the LZ2025 exclusion limit for spin-independent DM-nucleus scattering improves upon that of XENONnT by more than an order of magnitude. These updated bounds are not included in the present analysis.
@@ IV.3.3 Early universe constraints
CMB spectral distortions also test the nature of the dark sector. These distortions, which arise from non-standard energy injection (e.g. from particle decays or annihilations) that disrupts the thermodynamic equilibrium between photons and free electrons after BBN ($z\lesssim 10^{8}$), can be probed and thus used as constraints on models [390]. The FIRAS instrument on COBE measured the CMB energy spectrum, and it found it consistent with a perfect black-body spectrum at $T_{0}=2.725\pm 0.002$ K [391, 392]. For mediator decays with lifetimes $10^{4}$ s $\tau<10^{13}$ s, CMB spectral distortions provide stronger constraints than CMB anisotropies, though they are still weaker than those originating from BBN. Future experiments such as PIXIE or PRISM have nevertheless the potential to significantly improve the bounds [390].
Big Bang Nucleosynthesis, which occurred at $T_{\rm BBN}\sim 0.1\,$MeV, is another critical probe of dark matter. New particles can affect the primordial abundances of light nuclei by altering the Hubble rate or the entropy density of the universe. This leads to constraints on $N_{\rm eff}$, which are comparable to those originating from the CMB [393], but less stringent than the Lyman-$\alpha$ forest bounds discussed below. In addition, if the mediator has a lifetime longer than approximately $0.1$ s, its decay can induce non-thermal nuclear reactions during or after BBN, modifying standard predictions [394, 395]. Such constraints are especially relevant for superWIMP production scenarios involving mediators with lifetimes $10^{4}$ s $\tau<10^{13}$ s that decay into hadronic final states. For example, in top-philic scenarios, these constraints help close the existing gap between constraints stemming from Lyman-$\alpha$ forest probes and the collider limits [201, 204].
@@
[…]
@@ IV.4.1.1 Minimal quark-philic models in the canonical freeze-out regime
Figure 42: Constraints on minimal simplified $t$-channel DM models from cosmological and astrophysical observables, as well as from the measured $Z$-boson visible decay width. The coloured region in the $(M_{X},M_{Y}/M_{X}-1)$ plane represents scenarios that achieve $\Omega h^{2}\simeq 0.12$, with the coupling value $\lambda$ indicated by the grey-scale colour map. The left (right) panels correspond to models with self-conjugate (complex) DM, featuring scalar (top row), fermionic (middle row), and vector (bottom row) DM particles. The hatched regions indicate exclusions from gamma-ray searches (‘ID gamma rays’, including gamma-line searches from Fermi-LAT [353] and HESS [403] from the galactic centre and gamma-ray continuum searches in dSPhs by Fermi-LAT [348]), cosmic-ray antiproton searches (‘ID antiprotons’), DM direct detection via spin-independent and spin-dependent interactions (‘$D$ SI’ and ‘$D$ SD,’ respectively, including limits from LZ [404], PICO-60 [377], CRESST-III [405] and DarkSide-50 [406]), and $Z$-boson visible decays (‘Z decay’, from [407]). The blank upper region corresponds to scenarios requiring non-perturbative couplings, while the white region at the bottom represents the CDFO regime (see section IV.4.1.2). This figure is taken from [25], where further details can be found.
@@
In the canonical freeze-out regime, direct detection constraints originating from spin-independent DM interactions with nuclei represent the strongest constraints across all six models. They are found to exclude the entire sampled parameter space for all three complex DM classes of scenarios, as shown in the right panels of figure 42. This leaves the conversion-driven freeze-out region as the sole viable regime for these models. In contrast, for self-conjugate DM models, portions of the parameter space remain unaffected by direct detection bounds. Nevertheless, combining constraints from direct detection (both spin-independent and spin-dependent bounds), indirect detection via gamma-ray and cosmic-ray antiproton observations, and robust limits from $Z$-boson visible decay measurements excludes significant portions of the parameter space, as depicted in the left panels of figure 42. For real scalar DM, these constraints exclude all scenarios with DM masses below $800\,\text{GeV}$ or mediator masses below $2\,\text{TeV}$. For Majorana DM and real vector DM, a similar exclusion pattern arises, except for isolated allowed regions featuring scenarios with $M_{X}\lesssim 4\,\text{GeV}$ and $100\,\text{GeV}\lesssim M_{X}\lesssim 200\,\text{GeV}$ respectively. In the latter cases, cosmic-ray antiproton data provide key constraints for $200\,\text{GeV}\lesssim M_{X}\lesssim 800\,\text{GeV}$, similar to the scalar DM case. Here, indirect detection bounds are particularly stringent for vector and Dirac DM, due to the S-wave nature of the annihilation process into quark pairs. For additional results on these models, we also refer to [339].
Figure 43: Cosmologically viable region of the parameter space of models featuring a DM state coupling to the right-handed top quark $t_{R}$ (with $\Omega h^{2}=0.12$), and constraints shown as functions of the DM mass $M_{X}$ and the relative mass splitting $M_{Y}/M_{X}-1$. The left panel corresponds to Majorana DM, while the right panel depicts a scenario with real scalar DM. These plots are based on the results from [201] and [27], respectively, and the direction detection ($D$) constraints on both plots are derived from the XENON1T bounds. The thick black line separates the WIMP region (above) from the CDFO region (see section IV.4.1.2), and the shaded regions denote exclusions due to the various experimental and theoretical constraints discussed in the text.
@@ IV.4.2.2 Frustrated dark matter
By construction, $2\to 2$ interactions between the DM and the SM occur only at one-loop order. Consequently, for most DM masses, DM annihilation is dominated by $XX$ scattering to at least four SM particles, which effectively reduce to $t$-channel XXXX annihilations to a pair of mediators when the mediators can be produced nearly on-shell. For perturbative Yukawa-like DM-mediator couplings $\lambda_{X}\lesssim\sqrt{4\pi}$, efficient DM annihilation requires these $t$-channel processes to be kinematically accessible; otherwise, DM tends to be overabundant under the assumption of a standard cosmological history. However, the fDM framework allows for a wide range of mediator masses and SM couplings, enabling the construction of numerous specific realisations with large experimentally viable regions in the model parameter space. The regions of the fDM parameter space that can produce a viable thermal relic are, in principle, testable via indirect searches for DM annihilation in the cosmos. While loop-suppressed $2\to 2$ annihilation processes leading to photons or a $\gamma Z$ final state are possible, the dominant experimental signature of the models is typically an apparent excess in the continuum spectra of gamma or cosmic rays originating from DM annihilation into quarks and/or charged leptons. For fDM scenarios at or below the TeV scale, the strongest constraints thus come from Fermi-LAT searches for gamma-ray production in nearby dwarf spheroidal galaxies. On the other hand, direct detection is fully loop-suppressed since there are no tree-level diagrams for fDM scattering off SM particles. Nonetheless, constraints from direct detection experiments such as XENON1T can be quite stringent. In scenarios where mediators carry weak hypercharge, spin-independent DM-nucleon scattering via off-shell photon exchange then provides the most robust constraints, that can also be interpreted as upper limits on the DM magnetic dipole moment [368]. In the absence of such processes, box diagrams involving mediators and SM fields can sometimes generate detectable direct detection signals [366] that may lead to competitive or even exceed limits from indirect detection. More conclusive quantitative statements require, however, specifying a particular fDM construction.
@@ IV.4.2.3 Composite dark matter
Figure 50: Constraints on the parameter space of a composite $t$-channel DM scenario featuring one DM state $X$ and one mediator state $Y$ that couples to the top quark with a strength $\lambda_{t}$. Results are given in the $(M_{X},M_{Y})$ plane, and the value of the coupling $\lambda_{t}$, given through the colour map, is determined in order to reproduce the observed relic density $\Omega_{CDM}h^{2}=0.1186$. The black lines correspond to isolines of constant coupling values with $\lambda_{t}=1$, $2$, 33, 44 and $5$.
@@ A.1 Description of the analysis
Bin 0 EM0: $200$ GeV $E_{T}^{\mathrm{miss}}<250$ GeV IM0: $E_{T}^{\mathrm{miss}}>200$ GeV
Bin 1 EM1: $250$ GeV $E_{T}^{\mathrm{miss}}<300$ GeV IM1: $E_{T}^{\mathrm{miss}}>250$ GeV
Bin 2 EM2: $300$ GeV $E_{T}^{\mathrm{miss}}<350$ GeV IM2: $E_{T}^{\mathrm{miss}}>300$ GeV
Bin 3 EM3: $350$ GeV $E_{T}^{\mathrm{miss}}<400$ GeV IM3: $E_{T}^{\mathrm{miss}}>350$ GeV
Bin 4 EM4: $400$ GeV $E_{T}^{\mathrm{miss}}<500$ GeV IM4: $E_{T}^{\mathrm{miss}}>400$ GeV
Bin 5 EM5: $500$ GeV $E_{T}^{\mathrm{miss}}<600$ GeV IM5: $E_{T}^{\mathrm{miss}}>500$ GeV
Bin 6 EM6: $600$ GeV $E_{T}^{\mathrm{miss}}<700$ GeV IM6: $E_{T}^{\mathrm{miss}}>600$ GeV
Bin 7 EM7: $700$ GeV $E_{T}^{\mathrm{miss}}<800$ GeV IM7: $E_{T}^{\mathrm{miss}}>700$ GeV
Bin 8 EM8: $800$ GeV $E_{T}^{\mathrm{miss}}<900$ GeV IM8: $E_{T}^{\mathrm{miss}}>800$ GeV
Bin 9 EM9: $900$ GeV $E_{T}^{\mathrm{miss}}<1000$ GeV IM9: $E_{T}^{\mathrm{miss}}>900$ GeV
Bin 10 EM10: $1000$ GeV $E_{T}^{\mathrm{miss}}<1100$ GeV IM10: $E_{T}^{\mathrm{miss}}>1000$ GeV
Bin 11 EM11: $1100$ GeV $E_{T}^{\mathrm{miss}}<1200$ GeV IM11: $E_{T}^{\mathrm{miss}}>1100$ GeV